OpenStudy (anonymous):

Find the inverse of f(x)=(1+3x)/(5-2x) Please show work!

OpenStudy (anonymous):

All you gotta do is switch the x and y and solve bro

OpenStudy (anonymous):

solve for x becuase f(x) would be y

OpenStudy (anonymous):

you want the steps?

OpenStudy (anonymous):

yeah, but would you solve for y? I got up to this: \[(5-2y)x=1+3y\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[(5-2y)x=1+3y\] \[5x-2xy=1+3y\] \[5x-1=3y-2xy\] \[5x-1=(3-2x)y\] \[\frac{5x-1}{3-2x}=y\]

OpenStudy (anonymous):

i just did this problem a few posts back, but i don't know how to retrieve it

OpenStudy (anonymous):

think of the inverse like the opposite...that's why you have to switch them. If you can, graph it and you'll be able to see it a lot easier.

OpenStudy (anonymous):

so \[f^{-1}(x)=\frac{5x-1}{3-2x}\] assuming i did the algebra right. it was the steps you wanted yes?

OpenStudy (anonymous):

@Mandolino, I deleted the question because I awarded you a medal but I still needed help. Sorry.

OpenStudy (anonymous):

@satellite73 Thanks!

OpenStudy (anonymous):

i think there is a sign error, if you move -2xy to the rhs then is should be positive. i got earlier (hope its right)\[y=\frac{5x-3}{2x+3}\]

OpenStudy (anonymous):

i think there is a sign error, if you move -2xy to the rhs then is should be positive. i got earlier (hope its right)\[y=\frac{5x-3}{2x+3}\]

OpenStudy (anonymous):

should be 5x-1 in the numerator